OpenStudy (anonymous):

y = x2 + 6x − 16, I need th vertex, I know that it is -3, -25 but i keep getting 43 for y.

jimthompson5910 (jim_thompson5910):

-3 is the correct x coordinate, so you plug this into the given equation to get y = x^2 + 6x - 16 y = (-3)^2 + 6(-3) - 16 ... plug in x = -3 y = 9 + 6(-3) - 16 ... note: (-3)^2 is positive 9 and NOT negative 9 y = 9 - 18 - 16 y = -25 So when x = -3, y is y = -25 This means the vertex is (-3, -25)

OpenStudy (anonymous):

ooooooooh! ok cool thanks

jimthompson5910 (jim_thompson5910):

yw